how do i find the missing sides of each of these right triangles?
please explain..
a=?,b=24,c=26
a=5,b=6,c=?
a=7,b=?.c=14

Answer by Edwin McCravy(20055)

Use the appropriate one of these three Pythagorean theorems:

1.  a = Öc² - b²

2.  b = Öc² - a²

3.  c = Öa² + b²

-------------------

a=?,b=24,c=26
We need a, so we use the first Pythagorean theorem above:

a = Öc² - b²

a = Ö26² - 24² = Ö676 - 576² = Ö100 = 10

-------------------

a=5,b=6,c=?
We need c, so we use the third Pythagorean theorem above:

c = Öa² + b²

c = Ö5² + 6² = Ö25 + 36 = Ö61 

-------------------

a=7,b=?.c=14
We need b, so we use the second Pythagorean theorem above:

b = Öc² - a²

b = Ö14² - 7² = Ö196 - 49 = Ö147 = Ö49*3 = 7Ö3

-------------------
Edwin